# Uncertainty in latent representations of variational autoencoders optimized for visual tasks

### v2

@@ Methods
Through this procedure, uncertainty curves were computed for all the combinations of morphing from labels 00 to $9$ (Figure S4b color curves).
To evaluate if a certain curve has the expected behavior of being maximally uncertain at intermediate $\lambda$ values, we fitted a second order polynomial function (Figure S4c gray curves) and checked whether it’s curvature was negative and the maximum was located inside a window of size $L$ centered respect to $\lambda=0.5$, for different $L$ values (Figure S4c).
Image corruption
The image corruption experiment was done over the complete test set. Each image was synthetically corrupted by applying a two-dimensional Gaussian kernel of covariance $\eta_{b}I_{2x2}$. The resulting blurred image was encoded into the models’ latent space, where the measure of uncertainty was computed. We compute the uncertainty of the network for a given input by Equation 10.
This process was repeated for all images, and the average uncertainty was computed. This was repeated for different $\eta_{b}$ values until the images were completely blurred out ($\eta_{b}=16$ equal to half the pixel width of the images) We repeated the experiment with pixel-noise corruption. In this case, each pixel was intensity corrupted by additive white Gaussian noise of null mean and variance $\eta_{p}^{2}$. The whole perturbed image was then rescaled in intensity to preserve the original range of intensities.$\eta_{p}$ values between 0 and 1 were considered, going from no perturbation to a strongly perturbed image.
Out of distribution experiments
@@
To study the capacity of the trained models to associate near-to-prior uncertainty for images from previously unseen distributions, we compared the latent representations of images in and out of the distribution used for training. The reconstruction quality was analysed by projecting the posterior mean back to the pixel-space using the decoder to observe how OOD inputs were represented. Latent uncertainty to a given observation was computed the same way as in the two previous experiments. We looked at models trained with all datasets (MNIST, ChestMNIST, cMNIST, cFashionMNIST, cChestMNIST and Van Hateren) and tested with the complementary domains and corresponding pixel-shuffled dataset (Figure 2c and Figure S5). In all cases, images from the OOD datasets were resized to coincide the size of images from the training set, and rescaled in pixel intensity to match the in-distribution pixel intensity distribution. For the models trained on Van Hateren, that is population wise z-scoring the OOD datasets. For all other models trained on MNIST domains, that is rescaling the Van Hateren’s dataset pixel intensity distribution such that it’s mean $\pm$ 3 deviations fit the range [0,1].
@@
We found the same behavior as shown in the main text when training models with x-ray images and testing on natural images (Figure S5d, rightmost block). Yet when testing on digits and fashion images, both models seem to fare comparably, with reported uncertainties that do not approach the prior, although the reconstructed images do not collapse to within domain examples. When training with natural images and testing on the other (Figure S5e), the distribution of reported uncertainties, which originally had two modes for within distribution images corresponding to more informative (high contrast) and less informative (low contrast) images, reduces to a single mode mode for VAEs and EA-VAEs. The asymmetry in the results when original and target domains are inverted suggests a hierarchy in the richness of the feature space learned from the statistics of the images.
Predictive uncertainty
@@
Data availability
All datasets used in this paper are publicly available. Van Hateren dataset of natural image patches of 40x40 pixels is provided in Ref. [20]. MNIST and FashionMNIST datasets can be downloaded via Pytorch, and ChestMNIST dataset can be downloaded and imported using the MEDMNIST official repository.
Code availability
Implementation code and instructions for reproducing results are freely available on GitHub [40].
Acknowledgments
This work was supported by Argentina’s National Scientific and Technical Research Council (CONICET) (RE), which covered the salaries of all the authors, the the European Union project RRF-2.3.1-21-2022-00004 within the framework of the Artificial Intelligence National Laboratory (GO) and the National Research, Development and Innovation Office under grant agreement ADVANCED 150361 (GO). The authors gratefully acknowledge NVIDIA Corporation for the GPU computing, and the support of Universidad Nacional del Litoral (Grants CAID-PIC-50220140100084LI, 50620190100145LI), ANPCyT (PICT-PRH-2019-00009, 2022-03-00593).
Supplementary Information
Fig. S1: a, Trained VAE (left) and EA-VAE (right) models for inference on natural images. The encoder’s input of size $M=1600$ is followed by two fully connected hidden layers with softplus activation function. Two separate fully connected layers of $D=1800$ units encode the mean and variance of $q_{\phi}(\mathbf{z}|\mathbf{x})$. For the VAE, one fully connected layer decodes the latent sample into a reconstructed patch. For the EA-VAE model, the described VAE was modified such that the last component of the sample of the posterior distribution passes through function g() defining the effective $s$ sample from $q_{\psi}(s|\mathbf{x})$. The posterior sampled $s$ value is globally multiplied with the decoder’s output, resulting in the final reconstructed patch. b, Trained VAE (left) and EA-VAE (right) models for inference on natural images when scaling variable assumed from a Gamma distribution. In this case, the scaling variable is inferred from a separate encoder. $k=2$ was chosen as the smallest integer number that qualitatively achieved a similar distribution to the dataset’s $\sigma_{pix}$ distribution. Computationally, the choice of a fixed integer $k$, simplifies the sampling procedure, since one can always sample from the same standard Gamma function and then reparameterize by the scale parameter. This enables simple gradient computations across the sampling process. c, Trained VAE (left) and EA-VAE (right) models for inference on MNIST classic and contrast-augmented domains (digits, chest x-rays and fashion). For the VAE, the encoder’s input of size $M=32$ is followed by three convolutional hidden layers with relu activation function. Two separate fully connected layers of $D$ units encode the mean and variance of $q_{\phi}(\mathbf{z}|\mathbf{x})$. A fully connected and 3 deconvolutional layers decode the latent sample. Output from last layer passes trough h() function giving place to the reconstructed image. For the EA-VAE, the described VAE model in was modified such that separate fully connected layers encode mean and variance of a Normal distribution. Its sample passes through function g() defining the effective $s$ sample from $q_{\psi}(s|\mathbf{x})$. The posterior sampled $s$ value is globally multiplied with the decoder’s output before going through function g(), resulting in the final reconstructed image.

### v3

@@ 1.2 On the relationship between the scale variable and out-of-distribution inputs
2 Supplementary Methods and Experiments
2.1 Cost function and implementation for VAE and EA-VAE variants
2.1.1 Inference on natural images
Given the selected parametrization for prior, posterior and likelihood described in the main text Section˜2.2, Equation˜1 takes the form of
$\mathcal{L}_{VAE}(\mathbf{x},\psi,\phi)=\frac{1}{2\sigma_{obs}^{2}}\sum_{i=1}^{M}(\hat{\mathrm{x}}_{i}-\mathrm{x}_{i})^{2}+\frac{M}{2}\log{(2\pi\sigma_{obs}^{2})}$
$+\beta\sum_{j=1}^{D}\left(-1+|\mu_{j}(\mathbf{x})|-\log{(b_{j}(\mathbf{x})})+b_{j}(\mathbf{x})\exp{(-\frac{|\mu_{j}(\mathbf{x})|}{b_{j}(\mathbf{x})})}\right).$ (S25)
It is worth mentioning that in cases when the observation noise is assumed to be drawn from an uncorrelated Normal distribution, then Equation˜1 can be rewritten by global multiplication of $2\sigma_{obs}^{2}$ so that the $\beta$ hyperparameter absorbes the assumed observation noise [10], that is
$\beta^{\prime}=2\sigma_{obs}^{2}\beta.$ (S26)
Further normalization can as well turn the first term in Equation˜1 into the mean squared error (MSE). Selecting a larger $\beta$ as done in $\beta$-VAEs then corresponds to performing inference assuming an enlarged observation noise, fostering disentanglement by use of an uncorrelated prior.
In the case of the EA-VAE, when modeling the scaling variable as a softplus activated Laplace distribution, the loss function in Equation˜2 takes the following form
$\mathcal{L}_{EA-VAE}(\mathbf{x},\psi,\phi,\chi)=\frac{1}{2\sigma_{obs}^{2}}\sum_{i=1}^{M}(\hat{\mathrm{x}}_{i}-\mathrm{x}_{i})^{2}+\frac{M}{2}\log{(2\pi\sigma_{obs}^{2})}$
$+\beta_{1}\sum_{j=1}^{D-1}\left(-1+|\mu_{j}(\mathbf{x})|-\log{(b_{j}(\mathbf{x})})+b_{j}(\mathbf{x})\exp{(-\frac{|\mu_{j}(\mathbf{x})|}{b_{j}(\mathbf{x})})}\right)$
$+\beta_{2}\left(-1+|\mu_{s}(\mathbf{x})|-\log{(b_{s}(\mathbf{x}))}+b_{s}(\mathbf{x})\exp{(-\frac{|\mu_{s}(\mathbf{x})|}{b_{s}(\mathbf{x})})}\right),$ (S27)
where $\mu_{s}$ and $b_{s}$ are the inferred parameters of the scaling variable posterior. Importantly, to derive Equation˜S27 it was taken into account that any invertible transformation applied to $s$ ensures that the Kullback-Leibler (KL) divergence in the cost function remains unaffected[5]. We note that the resulting distribution (Laplace transformed by Softplus activation) does not have tractable moments. Thus we approximated the true posterior mean with samples, given that the inferred variance of $s$ was low except below the observation noise, where slightly different samples become similarly close to 0 due to softplus activation (Figure˜S3). Also note the substitution of $D$ in the first term of Equation˜S25 for D1D-1 in Equation˜S27, since we keep the overall number of latent variables unchanged for a fair comparison.
When modeling the scaling variable as a Log-normal distribution, we trained a version of the EA-VAE with a $\textrm{Log-normal}(0,1)$ prior for $s$. Similar to the homogeneous encoder, we use a joint encoder. This encoder produces mean and scale for Laplace variational posterior and a mean and variance of a Normal for the lognormal scaling variable, which is exponentiated after sampling from a Normal distribution (see Figure˜S1a for detailed architecture). Same as the main text model, $\beta_{2}$ was decreased linearly from 10 to 1 in the course of the first 100 epochs. Here, the loss function of the model resembles Equation˜S27, with the second KL term replaced with the respective KL for Normal distributions
$\mathcal{L}_{EA-VAE}(\mathbf{x},\psi,\phi,\chi)=\frac{1}{2\sigma_{obs}^{2}}\sum_{i=1}^{M}(\hat{\mathrm{x}}_{i}-\mathrm{x}_{i})^{2}+\frac{M}{2}\log{(2\pi\sigma_{obs}^{2})}$
$+\beta_{1}\sum_{j=1}^{D-1}\left(-1+|\mu_{j}(\mathbf{x})|-\log{(b_{j}(\mathbf{x}))}+b_{j}(\mathbf{x})\exp{(-\frac{|\mu_{j}(\mathbf{x})|}{b_{j}(\mathbf{x})})}\right)$
$+\beta_{2}\frac{1}{2}\left(-1+(\nu(\mathbf{x}))^{2}-\log{(\xi^{2}(\mathbf{x}))}+\xi^{2}(\mathbf{x})\right),$ (S28)
where $\nu$ and $\xi^{2}$ are the mean and variance of a Normal distribution for the lognormal scaling variable.
The third variant modeled the scaling variable as a Gamma distribution, which naturally restricts the scaling variable to positive values. The Gamma distribution can be parametrized in terms of the shape and scale parameters $k$ and $\theta$ respectively. For simplicity, we assumed that posteriors $q_{\chi}(s|\mathbf{x})=\textrm{Gamma}(s;k,\theta(\mathbf{x}))$ have a fixed shape parameter $k=2$. Since $k$ is fixed, the posterior’s mean, mode and variance depend only on the inferred scale parameter $\theta$. For the prior distribution, the shape parameter was also fixed at $k=2$ and the scale parameter $\theta=1/\sqrt{2}$ was chosen such that the prior had unit variance $k\theta^{2}=1$. A constant shape parameter $k$ means that the posterior standard deviation is always proportional to the posterior mean. This choice is in line with Weber’s Law of perception [13], where the estimation error of a quantity is a constant fraction of its mean (see Figure˜S1b for detailed architecture). In this case, Equation˜2 can be expressed similar to Equation˜S27, with the second KL term based upon Gamma distributions instead of Laplace [8]
$\mathcal{L}_{EA-VAE}(\mathbf{x},\psi,\phi,\chi)=\frac{1}{2\sigma_{obs}^{2}}\sum_{i=1}^{M}(\hat{\mathrm{x}}_{i}-\mathrm{x}_{i})^{2}$
$+\beta_{1}\sum_{j=1}^{D-1}\left(-1+|\mu_{j}(\mathbf{x})|-\log{(b_{j}(\mathbf{x}))}+b_{j}(\mathbf{x})\exp{(-\frac{|\mu_{j}(\mathbf{x})|}{b_{j}(\mathbf{x})})}\right)$
$+\beta_{2}\,2(-\log\sqrt{2}-\log\theta(\mathbf{x})+\sqrt{2}\,\theta(\mathbf{x})-1).$ (S29)
Note as well that because of the normalization done to the Van Hateren dateset in this case, observation noise $\sigma_{obs}$ gets rescaled and reduced, which we absorb into the $\beta_{1}^{\prime}$ and $\beta_{2}^{\prime}$ hyperparameters (Supplementary Table S1).
2.1.2 Inference in other classic computer vision domains
The cost functions constructed for the MNIST, ChestMNIST and NIH-ChestXray14 models under the assumptions described in the main text Section˜2.3, were
$\mathcal{L}_{VAE}(\mathbf{x},\psi,\phi)=\sum_{i=1}^{M}-(\mathrm{x}_{i}\log(\hat{\mathrm{x}}_{i})+(1-\mathrm{x}_{i})\log(1-\hat{\mathrm{x}}_{i}))$
$+\beta^{\prime}\sum_{j=1}^{D}\frac{1}{2}\left(-1+(\mu_{j}(\mathbf{x}))^{2}-\log{(\sigma_{j}^{2}(\mathbf{x}))}+\sigma_{j}^{2}(\mathbf{x})\right),$ (S30)
$\mathcal{L}_{EA-VAE}(\mathbf{x},\psi,\phi,\chi)=\sum_{i=1}^{M}-(\mathrm{x}_{i}\log(\hat{\mathrm{x}}_{i})+(1-\mathrm{x}_{i})\log(1-\hat{\mathrm{x}}_{i}))$
$+\beta^{\prime}_{1}\sum_{j=1}^{D-1}\frac{1}{2}\left(-1+(\mu_{j}(\mathbf{x}))^{2}-\log{(\sigma_{j}^{2}(\mathbf{x}))}+\sigma_{j}^{2}(\mathbf{x})\right)$
$+\beta^{\prime}_{2}\frac{1}{2}\left(-1+(\nu(\mathbf{x}))^{2}-\log{(\xi^{2}(\mathbf{x})}+\xi^{2}(\mathbf{x})\right),$ (S31)
2.2 Post-training evaluation of VAEs and EA-VAEs for natural images
2.2.1 Latent receptive fields
For fully connected models, as employed for the natural image dataset, the receptive field $\mathbf{RF}_{j}$ associated to the $j$-th latent dimension was computed through the Spike Triggered Average (STA) technique on the test dataset for all j=1,,1800j=1,...,1800. The STA was estimated as
$\mathbf{RF}_{j}=\underset{\left\{\mathbf{x}\right\}}{\textrm{mean}}\,\mu_{j}(\mathbf{x})\,\mathbf{x},$ (S32)
where $\mu_{j}(\mathbf{x})$ is the $j$-th component of the mean of $\mathbf{x}$’s inferred posterior $q_{\phi}(\mathbf{z}|\mathbf{x})$.
Latent units were discerned as informative or non-informative according to their contribution to the dataset reconstruction. When computing the relative change in the images reconstructed when all units are active, or when a specific unit is replaced by its mean activation throughout the ensemble, it was observed that the units associated to orientation-sensitive filters are clustered with the most contribution, while the second group barely contribute. All results shown in the main text were computed using only the informative latent dimensions.
2.2.2 Uncertainty representation in the latent space
For this, given an ensemble of patches $\left\{\mathbf{x}\right\}$, we infer the contrast $s^{*}(\mathbf{x})$ for each patch as the across-pixel standard deviation of intensities: $s^{*}(\mathbf{x})=\sigma_{pix}(\mathbf{x})$.
The signal mean is defined as the average distance of the posterior’s mean to the prior’s mean (0 in this case) over all patches that have the same inferred contrast. Signal variance is computed as the variance of the posterior’s mean of patches of same contrast, averaged across latent dimensions. Finally, noise variance is defined as the posterior’s variance vector $\boldsymbol{\sigma}^{2}(\mathbf{x})$ averaged across latent dimensions and observations of equal inferred contrast $s$.
$\text{SM}(s_{i})=\underset{\left\{\mathbf{x}|s^{*}(\mathbf{x})\in B_{i}\right\}}{\textrm{mean}}\left\|\boldsymbol{\mu}(\mathbf{x})\right\|$ (S33)
$\text{SV}(s_{i})=\underset{j}{\textrm{mean}}\underset{\left\{\mathbf{x}|s^{*}(\mathbf{x})\in B_{i}\right\}}{\textrm{Var}}\mu_{j}(\mathbf{x})$ (S34)
$\text{NV}(s_{i})=\underset{j}{\textrm{mean}}\underset{\left\{\mathbf{x}|s^{*}(\mathbf{x})\in B_{i}\right\}}{\textrm{mean}}\sigma_{j}^{2}(\mathbf{x})$ (S35)
The associated errors can be expressed as
$\epsilon_{{}_{\text{SM}}}(s_{i})=\frac{1}{\sqrt{N_{i}}}\sqrt{\underset{\left\{\mathbf{x}|s^{*}(\mathbf{x})\in B_{i}\right\}}{\textrm{Var}}\left\|\boldsymbol{\mu}(\mathbf{x})\right\|}$ (S36)
$\epsilon_{{}_{\text{SV}}}(s_{i})=\frac{1}{\sqrt{D}}\underset{j}{\textrm{mean}}\sqrt{\frac{2}{(N_{i}-1)}}\underset{\left\{\mathbf{x}|s^{*}(\mathbf{x})\in B_{i}\right\}}{\textrm{Var}}\mu_{j}(\mathbf{x})$ (S37)
$\epsilon_{{}_{\text{NV}}}(s_{i})=\frac{1}{\sqrt{N_{i}}\sqrt{D}}\sqrt{\underset{\left\{j\cup\mathbf{x}|s^{*}(\mathbf{x})\in B_{i}\right\}}{\textrm{Var}}\boldsymbol{\sigma}^{2}(\mathbf{x})}.$ (S38)
It is worth mentioning that the natural scenes dataset used in this work has patches with varied values of pixel intensity standard deviation $\sigma_{pix}(\mathbf{x})$. As patches with high contrast live far away from the origin in the M-dimensional pixel space (since $s_{pix}(\mathbf{x})=\sqrt{\frac{1}{M-1}\sum_{i=1}^{M}(x_{i}-\bar{x})^{2}}=\sqrt{\frac{1}{M-1}}\left\|\mathbf{x}\right\|$), it is expected that the network will have more difficulty exploring farther regions in the pixel space than the regions where low-contrast patches live. Therefore, in evaluation instances, patches with contrast higher than $\sigma_{pix}=2$ were not considered.
2.2.3 Divisive normalization experiments
To characterize the dependencies between latents of the Variational Autoencoders, we calculated conditional distributions for pairs of latents by analyzing the distribution of posterior means for a large number of natural image patches. Across a large number of randomly sampled filter pairs the variance of the conditional distribution was calculated at four quadrants: two central and two flanking quadrants (Figure˜7a), which corresponded to low-level, and high-level activations of the conditioned filter, respectively (Figure˜7b).
$z_{i}=\frac{L_{i}}{\sqrt{\sum_{j\neq i}w_{ji}L_{j}^{2}+\sigma^{2}}}.$ (S39)
2.3 Post-training evaluation of VAEs and EA-VAEs for MNIST, ChestMNIST and NIH-ChestXray14
2.3.1 Out of distribution experiments
@@
To study the capacity of the trained models to associate near-to-prior uncertainty for images from previously unseen distributions, we compared the latent representations of images in and out of the distribution used for training. The reconstruction quality was analysed by projecting the posterior mean back to the pixel-space using the decoder to observe how OOD inputs were represented. Latent uncertainty to a given observation was computed by Equation˜4. We looked at models trained with all datasets (MNIST, ChestMNIST, cMNIST, cFashionMNIST, cChestMNIST and Van Hateren) and tested with the complementary domains and corresponding pixel-shuffled dataset (Figure˜3a and Figure˜S5). In all cases, images from the OOD datasets were resized to coincide the size of images from the training set, and rescaled in pixel intensity to match the in-distribution pixel intensity distribution. For the models trained on Van Hateren, that is population wise z-scoring the OOD datasets. For all other models trained on MNIST domains, that is rescaling the Van Hateren’s dataset pixel intensity distribution such that it’s mean $\pm$ 3 deviations fit the range [0,1].
@@
The same findings as the ones described in the main text Section˜4.3 were replicated when training models with ChestMNIST and testing on natural images (Figure˜S5d, rightmost block), yet when testing on digits and fashion images, both models seem to fare comparably, with reported uncertainties that do not approach the prior, although the reconstructed images do not collapse to within domain examples. When training with natural images and testing on the other (Figure˜S5e), the distribution of reported uncertainties, which originally had two modes for within distribution images corresponding to more informative (high contrast) and less informative (low contrast) images, reduces to a single mode mode for VAEs and EA-VAEs. The asymmetry in the results when original and target domains are inverted suggests a hierarchy in the richness of the feature space learned from the statistics of the images.
2.3.2 Image corruption
The image corruption experiment was done over the complete test set. Each image was synthetically corrupted by applying a two-dimensional Gaussian kernel of covariance $\eta_{b}I_{2x2}$. The resulting blurred image was encoded into the models’ latent space, where the measure of uncertainty was computed. We compute the uncertainty of the network for a given input by Equation˜4. This process was repeated for all images, and the average uncertainty was computed. This was repeated for different $\eta_{b}$ values until the images were completely blurred out ($\eta_{b}=100$ in the case of NIH-ChestXray14 dataset). We repeated the experiment with pixel-noise corruption. In this case, each pixel was intensity corrupted by additive white Gaussian noise of null mean and variance $\eta_{p}^{2}$. The whole perturbed image was then rescaled in intensity to preserve the original range of intensities. $\eta_{p}$ values between 0 and 2.5 were considered, going from no perturbation to a strongly perturbed image.
2.3.3 Uninformative observations
We evaluated the models’ ability to handle uninformative observations by creating an “average image” through pixel-wise averaging of all dataset images. This image served as a proxy for an uninformative observation. As expected, this image resembles a blurry, unrecognizable digit (MNIST) or a slightly blurry x-ray (ChestMNIST) (Figure˜S8). We characterize the latent uncertainty in the models measuring the noise standard deviation of the posterior and averaging along the $\mathbf{z}$ latent dimensions (Equation˜4). In the case of MNIST, the VAE produced a latent uncertainty ($u_{{}_{\text{VAE}}}=0.17$) comparable to that of real handwritten digits ($u_{{}_{\text{ID}}}$) in the dataset $p(u_{{}_{\text{ID}}}>u_{{}_{\text{VAE}}})=0.49$], and considerably below the unit standard deviation of the prior (Figure˜S8 left panel). Notably, the same measurement in the EA-VAE yielded $u_{{}_{\text{EA-VAE}}}=0.36$, This uncertainty is not only twice of that obtained for the VAE but also significantly higher than the uncertainty of actual samples from the distribution $p(u_{{}_{\text{ID}}}>u_{{}_{\text{EA-VAE}}})=10^{-4}$]. A tendency similar to that observed on the MNIST dataset can also be seen on ChestMNIST-trained VAE and EA-VAE models $p(u_{{}_{\text{ID}}}>u_{{}_{\text{VAE}}})=0.33$; $p(u_{{}_{\text{ID}}}>u_{{}_{\text{EA-VAE}}})=0.12$] (Figure˜S8 right panel). Note, that the average x-ray image resembles actual data samples more closely than the average digit image, which explains the less pronounced advantage of EA-VAE over the standard VAE. In summary, these results show that uncertainty estimates in the EA-VAE, but not in the standard VAE, distinguish uninformative images from actual images.
2.3.4 Morphing and uncertainty
We conducted an experiment to study the uncertainty in the VAE and EA-VAE latent space when morphing digits from one label to another. The first step for doing this was selecting all the test images that belong to a specific label A and computing the average coordinate in the latent space of these (Figure˜S6a bottom panels). The output $\hat{\mathbf{x}}_{A}$ of the decoder from that position is a representative image of the label A (Figure˜S6a top). The same procedure is repeated with images belonging to label B obtaining $\hat{\mathbf{x}}_{B}$. Then, morphing is done by linearly combining $\hat{\mathbf{x}}_{A}$ and $\hat{\mathbf{x}}_{B}$ with weight $\lambda$ varying from 0 to 1, that is
$\hat{\mathbf{x}}_{AB,\lambda}=(1-\lambda)\hat{\mathbf{x}}_{A}+\lambda\hat{\mathbf{x}}_{B}.$ (S40)
We then inferred the corresponding posteriors taking these morphed images as input, and computed the uncertainty of the network by Equation˜4.
Through this procedure, uncertainty curves were computed for all the combinations of morphing from labels 0 to $9$ (Figure˜S6b color curves).
To evaluate if a certain curve has the expected behavior of being maximally uncertain at intermediate $\lambda$ values, we fitted a second order polynomial function (Figure˜S6c gray curves) and checked whether it’s curvature was negative and the maximum was located inside a window of size $L$ centered respect to $\lambda=0.5$, for different $L$ values (Figure˜S6c).
2.3.5 Predictive uncertainty
@@ 2.3.6 MNIST MLP classifier
2.4 Uncertainty Representations for ImageNet in Inception Feature Space
Inception-v3 [11] has become a canonical architecture for computer vision. The activations of this model, when pretrained with Imagenet [3, 9], are nowadays often employed to construct a representational space where to compare distributions of images: for instance between real and synthetically generated data [6].
In this section, we compare the capabilities of VAEs and EA-VAEs to develop a meaningful representation of uncertainty, and detect distributional shifts in the Inception representational space. To that end, we first precompute Inception-v3 features for images from ImageNet (Figure˜S10a). We employ the three categories of human images present in Imagenet: ‘Ball Player’, ‘Groom’ and ‘Scuba Diver’, and select another three categories: ‘Desk’, ‘Wall Clock’ and ‘Bookcase’ (Figure˜S10b). To analyze whether VAEs and EA-VAEs can learn meaningful representations of uncertainty that capture the semantic similarity between classes, we train these models on the first of the human classes (‘Ball Player’), and evaluate reported uncertainty on examples from the other human and object classes (Figure˜S10c). One would expect the models to be able to detect out of distribution examples, with higher uncertainty for those classes which are more dissimilar to the training set.
To systematically compare VAE and EA-VAE uncertainty representations for out of distribution images, we trained $N=10$ VAEs and EA-VAEs. Encoders and decoders are symmetric with two fully connected layers each (of 512 and 256 neurons with relu activation), representing posteriors with 64 latent dimensions. We kept the VAE and EA-VAE architectures as similar as possible to each other. The only difference being that in the EA-VAE the last of the 64 latents went through a soft-plus function before multiplying the output, instead of integrating the standard decoder as the rest of the variables. Otherwise, the cost function remains intact, and all latent variables have identical normal priors. For each model, we obtained the distribution of posterior latent uncertainties for each image class (Figure˜S10c), and then we computed the p-values for a t-test between the distribution of posterior uncertainties for out-of-distribution examples vs in-distribution examples for both models (Figure˜S10d). We observe that the EA-VAE shows consistently higher significance than the VAE in the t-test for object images. This tendency is reversed for the ‘Groom’ category, indicating that the EA-VAE perceives these two human classes as more similar than the VAE. Initially one may have expected the ‘Scuba Diver’ class to be regarded as more similar by these models, since the contain humans. We observe however that both models report very high uncertainty in this case, surpassing that of objects. Close inspection of the images (see third column in Figure˜S10b) shows that these images are indeed quite dissimilar to the rest, with underwater scenes, and a major shift in the distribution of colors and shapes.
2.5 Uncertainty Representations for Importance Weighted Autoencoder
Importance Weighted Autoencoders (IWAEs) are a flexible extension of the VAE family, enabling a richer latent representation by taking multiple samples from the posterior [1]. This family of models optimizes the standard ELBO while permitting a more complex implicit posterior distribution that loosens the usual independence assumptions on the latent variables [2]. IWAEs could be principled candidates for modeling natural images, where latents have been shown to display strong nonlinear dependencies [12].
IWAEs are trained by optimizing for the following objective, which is equivalent to the VAE ELBO in expectation:
$\mathcal{L}_{\mathrm{IWAE_{k}}}(\mathbf{x},\psi,\phi)=\mathbb{E}_{\mathbf{z_{1}},\ldots,\mathbf{z_{k}}\sim q_{\phi}(\mathbf{z}\mid\mathbf{x})}\left[\log\left(\frac{1}{k}\sum_{i=1}^{k}\frac{p_{\psi}(\mathbf{x}\mid\mathbf{z_{i}})p(\mathbf{z_{i}})}{q_{\phi}(\mathbf{z_{i}}\mid\mathbf{x})}\right)\right]$ (S41)
Here, the encoder and decoder are parameterized by $\phi$ and $\psi$, respectively and $k$ is a fixed hyperparameter of the model. The output of the encoder, $q_{\phi}(\mathbf{z}\mid\mathbf{x})$ is the naive posterior obtained by the VAE independence assumptions. Drawing $k$ samples from the naive posterior and implementing an importance weighting over them leads to a richer posterior distribution. This does not take an explicit form, but it is possible to draw samples from this implicit distribution.
We trained IWAEs on the Van Hateren dataset to see whether this extension leads to remedied uncertainty representation. The models were trained on 40x40 pixel patches with the same settings as the standard VAE. Three models were tested, using $k=5,10,50$ samples from the naive posterior distribution. For the analysis, we took 15 samples from the posterior and calculated the sample mean and sample variance for estimates of the implicit posterior moments. The results were no different than in the case of the standard VAE: posterior variance still fails to capture the uncertainty related to low-contrast images (Figure˜S11). This shows that even a more flexible variant of the VAE is unable to perform a fundamental task of uncertainty estimation, which is remedied by the EA-VAE model.
2.6 Uncertainty Representations for Monte Carlo Dropout VAE
A common approach to approximate uncertainty in deep neural networks is Monte Carlo dropout [4], by maintaining dropout active during inference time and performing multiple stochastic forward passes[7]. To assess whether the EA-VAE achieves improved uncertainty representations beyond standard baselines, we compared it as well with a VAE with Monte Carlo dropout.
Following common practice in Bayesian deep learning, stochasticity (via dropout layers) was introduced exclusively in the encoder network, which models the approximate posterior over latent variables. The decoder was kept deterministic given the latent sample. During training, dropout was applied in the standard way, and during inference time, we kept dropout explicitly active.
At inference, for a given input image $\mathbf{x}$, the approximate posterior $\tilde{q}(\mathbf{z}|\mathbf{x})\approx\textrm{Normal}(\mathbf{z};\tilde{\boldsymbol{\mu}}(\mathbf{x}),\tilde{\boldsymbol{\sigma}}^{2}(\mathbf{x}))$ is then estimated through $T$ forward passes of the trained encoder, obtaining a set of $T$ latent posterior parameters from $q_{\phi}(\mathbf{z}|\mathbf{x})=\textrm{Normal}(\mathbf{z};\boldsymbol{\mu}(\mathbf{x}),\boldsymbol{\sigma}^{2}(\mathbf{x}))$
$\left\{\boldsymbol{\mu}(\mathbf{x})_{t},\boldsymbol{\sigma}^{2}(\mathbf{x})_{t}\right\}_{t=1}^{T},$ (S42)
where each pair corresponds to one realization of the dropout mask.
The aggregated posterior mean $\tilde{\boldsymbol{\mu}}(\mathbf{x})$ and variance $\tilde{\boldsymbol{\sigma}}^{2}(\mathbf{x})$ for a single image $\mathbf{x}$ were then computed using the law of total expectation and law of total variance
$\tilde{\boldsymbol{\mu}}(\mathbf{x})=\mathbb{E}_{t}[\boldsymbol{\mu}(\mathbf{x})_{t}]$ (S43)
$\tilde{\boldsymbol{\sigma}}^{2}(\mathbf{x})=\mathbb{E}_{t}[\boldsymbol{\sigma}^{2}(\mathbf{x})_{t}]+\text{Var}_{t}[\boldsymbol{\mu}(\mathbf{x})_{t}]$ (S44)
The VAE with Monte Carlo dropout was trained on the manuscript digits MNIST dataset, following the same architecture and hyperparameters used for the MNIST-trained standard VAE and EA-VAE (Figure˜S1c and Table Table˜S2). The model was trained with dropout probability hyperparameter set as $p=0.2$ and $p=0.5$, without observing qualitative differences in the results.
To evaluate the model in its capacity to represent uncertainty faithfully, we reproduced the experiments done for the standard VAE and EA-VAE on uninformative observations (main text Section˜2.3.3 and Figure˜S12a), morphing digits (main text Section˜4.5 and Figure˜S12b) and out of distribution detection (main text Section˜4.3 and Figure˜S12c). In all cases, the VAE with Monte Carlo dropout failed to represent uncertainty in a meaningful way, equivalently to how the Standard VAE behaved.
3 Supplementary Tables and Figures
Table S1: Hyperparameter settings for the Van Hateren models